I have the following word problem and really do not know where to begin figuring it out, using algebra anyways. Please Help!!
Allen bought 20 stamps at the post office in 37¢ and 20¢ denominations. If the total cost of the stamps was $7.06, how many 37¢ stamps did Allen buy?

Answer by stanbon(75887)

Allen bought 20 stamps at the post office in 37¢ and 20¢ denominations. If the total cost of the stamps was $7.06, how many 37¢ stamps did Allen buy?
-------------
Let number of 37 cent stamps be "x"; value of those is 37x cents
The number of 20 cent stamps is "20-x" ;
value of those is 20(20-x)=400-20x cents
-----------
EQUATION:
value + value = 706 cents
37x + 400-20x = 706
17x = 306
x= 18 (This is number of 37 cent stamps)
20-x = 2 (This is number of 20 cent stamps)
Cheers,
Stan H.